Arteriosclerotic plaque narrows down a section of an artery
to 20% of its normal cross-sectional area. What is the
pressure in that section if immediately before it the pressure
is 100 mmHg and the flow speed is 0.12 m/s?
1
Expert's answer
2019-09-27T09:44:59-0400

Equation of continuity:


"A_1v_1=A_2v_2"

"A_1v_1=0.2A_1v_2"

The flow speed in this section:


"v_2=5v_1=5(0.12)=0.6\frac{m}{s}"

Bernoulli’s Equation:


"p_1+0.5\rho v_1^2=p_2+0.5\rho v_2^2"

"13332+0.5(1050)(0.12)^2=p_2+0.5(1050)(0.6)^2"

"p_2=13151\ Pa"